OpenStudy (anonymous):

Solve the following system by Gauss–Jordan elimination. 2x₁+5x₂+3x₃=35 12x₁+31x₂+20x₃=216

OpenStudy (anonymous):

I got x₁=5/2 x₂=6 x₃=t (free variable) The answer says I am wrong. Care to tell me what I did wrong?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ParthKohli @UnkleRhaukus

OpenStudy (unklerhaukus):

x_3=0

OpenStudy (anonymous):

Right. But whats wrong with my other answers?

OpenStudy (anonymous):

Here is what the online page looks like:

OpenStudy (anonymous):

Shouldn't the other variables be in terms of the free variable?

OpenStudy (anonymous):

What did were your elimination steps?

OpenStudy (anonymous):

It's difficult to post the elimination steps here. I checked with my calculator though. It's correct.

OpenStudy (anonymous):

But yeah, that's what I think @wio . I don't know how to express it though.

OpenStudy (unklerhaukus):

when we check your solution , we see t must be 0 x₁=5/2 x₂=6 x₃=t _________ 2x₁+5x₂+3x₃=35 5+30+3t = 35 35 +3t = 35 _________ 12x₁+31x₂+20x₃ = 216 60 + 186 + 20t = 216 216 + 20t = 216

OpenStudy (anonymous):

but t is a free variable is it not? SO it can be any value it wants.

OpenStudy (anonymous):

\[ \begin{bmatrix} 2&5&3 \\ 12&31&20 \\ 0&0&1 \end{bmatrix} \begin{bmatrix} x_1 \\ x_2 \\ x_3 \end{bmatrix} = \begin{bmatrix} 35 \\ 216 \\ t \end{bmatrix} \]

OpenStudy (anonymous):

Did you solve it like this?

OpenStudy (anonymous):

No... How did you get 0,0 and 1?

OpenStudy (anonymous):

I just added in the equation \(x_3=t\)

OpenStudy (anonymous):

Sorry, that doesn't make sense.

OpenStudy (anonymous):

I added in the equation: \[ 0x_1+ 0x_2 +1x_3 = t \]

OpenStudy (anonymous):

Ohh I see.

OpenStudy (anonymous):

All of the elimination you did before should work for the most part.

OpenStudy (anonymous):

So I do a back substitution?

OpenStudy (anonymous):

You first want to get the second row to have a zero in it's first colum

OpenStudy (anonymous):

because x3=t so solve the others n terms of t.

OpenStudy (anonymous):

then you can back substitute

OpenStudy (anonymous):

1 second.

OpenStudy (anonymous):

x₁=2.5+3.5t ?

OpenStudy (anonymous):

x₂=6-2t?

OpenStudy (anonymous):

yeah I'm getting that too

OpenStudy (anonymous):

did you try it?

OpenStudy (anonymous):

One attempt left :3 . Fingers crossed.

OpenStudy (anonymous):

you can double check it you know

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

plug in each equation into the equations

OpenStudy (anonymous):

Assign a random number to t and see if it works?

OpenStudy (anonymous):

like, plug in t for x_3 and 6-2t for x_2 into the equations

OpenStudy (anonymous):

hmm actually I'm not quite sure how to double check it

OpenStudy (anonymous):

let me look at wolfram

OpenStudy (anonymous):

You can do this on wolfram? O_O .

OpenStudy (anonymous):

wolfram agrees... it's just a matter of the stupid answer verify thing not being a wingspan

OpenStudy (anonymous):

THANKS!!! It's right :) .

OpenStudy (anonymous):

No problem. Matrix is your friend.

OpenStudy (anonymous):

I love Linear algebra <3 .